OpenStudy (anonymous):

A 30.0 g arrow is shot by William Tell though an 8.00 cm thinck apple sitting on top of his son's head. If the arrow enters the apple at 30.0 m/s and emerges at 25.0 m/s in the same direction, with what force has the apple resisted the arrow?

OpenStudy (anonymous):

Lets consider the newtons 2nd law: F = ma Given: Mass , m = 30.0 g = 0.3 kg length of thickness , l = 8.00cm = 0.08m Accelaration , a = (vf^2 - vi^2 ) / 2 (0.08) = -1720 m/s^2 (Using third equation of motion) Force= m * a = -51.6 N but this force will be opposing and with a negative sign.

OpenStudy (anonymous):

Um... I didnt get quite the same numbers, could you work it out please?

OpenStudy (anonymous):

I have done raw calculations, better use a calculator ..the formula's are correct so you don't need to follow my answer ....I just gave the most accurate idea ..:)

OpenStudy (anonymous):

Okay, thank you ^^